The lab technician would like to prepare a 25 L solution of density 2 g/L. What mass of the solute is required for the solution?


1
Expert's answer
2021-02-04T07:05:42-0500

"\rho = \frac{m}{V}; m = \rho V;"

m = 2 * 25 = 50 (grams)


Solution: 50g.